OpenStudy (anonymous):

What are the solutions to the equation y² –5 = 31? Choose all answers that are correct. A. 18 B. 6 C. –18 D. –6

OpenStudy (anonymous):

It is B. All you do is take the numbers from the choices and plug them in, like 6^2 -5, if it equals 31 is it true. B does equal 31 so that is your answer,

OpenStudy (the_fizicx99):

Do the inverse of everything to isolate y! y^2 - 5 = 31 Add 5 to both sides, y^2 -5 = 31 +5 +5 y^2 = 36 Now sqaureROOT both sides, to undo the square, √y^2 = y √36 = 6 So y = 6

OpenStudy (anonymous):

its a multi awnser question

OpenStudy (anonymous):

B is the only correct one I can get out of it, you can also try tHe_FiZiCx99 way also, but I found mine to be easier.

OpenStudy (the_fizicx99):

That's actually considered cheating, if you didn't have the options you would be stuck o.o

OpenStudy (anonymous):

thank you i passed

OpenStudy (the_fizicx99):

Yw :>

OpenStudy (anonymous):

I know how to do it your way, I just don't if they give me the options. It's not cheating, that's how I get my answer. Everyone finds it differently.

OpenStudy (the_fizicx99):

My point is if you didn't have the options you'd be stuck if you didn't know how to solve it.

OpenStudy (anonymous):

I clearly just said that I do know how to solve it without the options.

OpenStudy (the_fizicx99):

I was restating what I said...

OpenStudy (anonymous):

Yeah, well I read it clearly the first time so no need to repeat yourself. I know what I am doing, I'm in high school. Thanks.

OpenStudy (anonymous):

O~O

OpenStudy (the_fizicx99):

^^ k. No need to argue. Bye!